The volume of a sphere is given by 4/3πr^3, find the volume of a hemisphere when the diameter is 48m.
If anyone could help me with this i would be very grateful as i have been having problems with it.
Thanks

Answer by stanbon(75887)

The volume of a sphere is given by 4/3πr^3, find the volume of a hemisphere when the diameter is 48m.
-----------------
A hemisphere is (1/2) of a sphere.
If the diameter is 48m, the radius is 24m.
---------------------
Vol. = (4/3)(pi)r^3
Your Problem:
Vol = (1/2)(4/3)(pi)(24)^3
Vol = (2/3)(pi)13824
Vol = 28952.92 cu.meters
=================
Cheers,
Stan H.